# Feferman’s completeness theorem

### v1

@@ 1. Introduction
By Gödel’s Second Incompleteness Theorem no consistent extension of $\mathsf{PA}$ with a computably enumerable axiomatization can prove its own consistency. Thus, whenever we extend a sound arithmetical theory $T$ to a sound theory $T^{\prime}$ that proves the consistency of $T$, it is guaranteed that $T^{\prime}$ is stronger than $T$, i.e., $Th(T^{\prime})\supseteq Th(T)$. The idea of using this phenomenon to attain transfinite sequences of theories of ascending strength is going back to Alan Turing [turing1939]. In modern terms, Turing defined transfinite iterations of local reflection $\mathsf{Rfn}$ by
@@
A subtle feature of this definition is that it requires to pick some fixed c.e. axiomatization of $\mathsf{Rfn}^{\alpha}(T)$ at each step. Thus $\alpha$ have to be computable ordinals with fixed computable presentations and the theories $\mathsf{Rfn}^{\alpha}(T)$ depend not only on the order type of $\alpha$ but also on the particular choice of presentation. As was proven already by Turing, for any true $\Pi_{1}$-sentence $\varphi$ there is a computable presentation $\alpha$ of the ordinal $\omega+1$, in fact an ordinal notation in Kleene’s $\mathcal{O}$, such that $\mathsf{Rfn}^{\alpha}(\mathsf{PA})$ proves $\varphi$. Of course there cannot be a single computable presentation $\alpha$ such that $\mathsf{Rfn}^{\alpha}(\mathsf{PA})$ proves every true $\Pi_{1}$ sentence, as this would mean that there is a c.e. extension of $\mathsf{PA}$ proving all true $\Pi_{1}$ statements. Thus, the construction of theories $\mathsf{Rfn}^{\alpha}(\mathsf{PA})$ has intensional character: Their consequences depend not only on the order type of $\alpha$, but also on the choice of particular computable presentation.
Turing also conjectured that for every arithmetical theorem $\varphi$, there is a computable presentation of an ordinal $\alpha$ such that $\mathsf{Rfn}^{\alpha}(\mathsf{PA})$ proves $\varphi$. Later, Feferman [feferman1962] refuted this conjecture and considered a variant of Turing’s progression, where local reflections $\mathsf{Rfn}(T)$ are replaced with uniform reflections $\mathsf{RFN}(T)$:
@@
For this variant of progressions, Feferman proved that Turing’s conjecture holds, i.e., any true arithmetical sentence is provable in $\mathsf{RFN}^{\alpha}(\mathsf{PA})$. This result is often referred to as Feferman’s completeness theorem. Feferman’s construction gives $\omega^{{\omega}^{\omega+1}}$ as an upper bound on the order type of the ordinal notations along which one needs to iterate reflection to obtain a prove of $\varphi$.
The presence of an upper bound in Feferman’s result indicates that it is a negative result: It essentially shows that proof-theoretic strength of an arithmetical sentence $\varphi$ couldn’t be appropriately measured by the least order type of ordinal notation, iteration of reflection along which proves $\varphi$. However, when restricted to natural ordinal notation systems the length of iterations of reflection necessary to obtain a particular theory does reflect the proof theoretic strength of the theory. This phenomenon was explored and used for the applications to ordinal analysis [schmerl1979fine, beklemishev2003proof]. We furthermore note that as was shown by the first author and James Walsh [pakhomov2021reflection], for systems of second-order arithmetic and their $\Pi^{1}_{1}$-consequences, the strength of the systems can be appropriately measured in terms of the length of iterations of $\Pi^{1}_{1}$-reflection, which is a measure closely connected with the $\Pi^{1}_{1}$-proof theoretic ordinals of theories.
Feferman’s original proof of his completeness theorem is not widely known; likely, due to its considerable technical difficulty. One feature that Feferman’s proof shares with Turing’s (rather simple) proof of the $\Pi^{0}_{1}$-completeness of consistency progressions is the employment of non-standard definitions of the axioms of theories. These non-standard definitions are engineered in such a way that we know them to be an axiomatization of a “natural” theory only if we know a certain sentence $\varphi$ to be true. In Feferman’s proof they appear in ever more intertwined ways. Very artfully, he applies increasingly more entangled versions of the recursion theorem to generate non-standard definitions of theories (see [franzen2004] for further details). Franzén [franzen2004, p. 387] expressed this as follows:
The completeness theorem [i.e. Feferman’s] can be seen as a dramatic illustration of the role of intensionality in logic.
In this paper we present three new proofs of Feferman’s completeness theorem. The first is a simple proof of this theorem using results not available to Feferman at the time such as the conservativity of $\mathsf{ACA_{0}}$ over $\mathsf{PA}$. The second proof is an alternative proof using techniques that Feferman could have had access to, such as Schütte’s completeness theorem for $\omega$-logic. At last we combine our first proof with results of Ash and Knight [ash1990] in computable structure theory to obtain precise bounds on the order types appearing in Feferman’s theorem.
@@ 1.1. Structure of the paper
    For any computable linear order $\mathcal{L}$ such that $\mathsf{ACA}_{0}\vdash\mathsf{LO}(\mathcal{L})$ the theories $\mathsf{ACA}_{0}+\mathsf{WO}(\mathcal{L})$ and $\mathsf{PA}+\mathsf{LO}(\mathcal{L})+\mathsf{TI}(\mathcal{L})$ prove exactly the same first-order sentences.
@@
The ideas of the proofs of this components are the following:
@@
    This is a particular case of $\mathsf{ACA}_{0}$-provable $\Pi^{1}_{1}$-completeness of well-orders. For the convenience of readers we give a standard proof of the fact (Lemma 1)
@@
    We prove this in Lemma 16 by showing that the extension of any model of $\mathsf{PA}+(\mathcal{L})+\mathsf{TI}(\mathcal{L})$ by the second-order universe consisting of first-order definable sets is a model of $\mathsf{ACA}_{0}+\mathsf{WO}(\mathcal{L})$.
@@
    Like most other facts about the iterations of reflection principles, this fact is straightforward to prove using Löb’s Theorem
We note that neither of this facts are actually new. The first is going back to Kleene and the formalized version can be found in [simpson2009]. The second fact is a triviality, while the third one is folklore.
@@
To prove Theorem 1.1 we replace the use of $\Pi^{1}_{1}$-completeness of computable well-orders with a sharper theorem of Ash and Knight [ash1990], who proved the $\Pi^{0}_{2n+1}$-hardness of the property to be an index of a computable order of order type $\omega^{n}$. [^1] In fact it is easy to see that this set of indices is $\Pi^{0}_{2n+1}$-complete
We complement Theorem 1.1 with the following result proving that the upper bound on order types is in fact sharp:
@@
theoremlowerbound There exists a true $\Pi_{2n}$-sentence $\varphi$ such that for any computable $\mathsf{PA}$-verifiable linear order $\mathcal{L}$ of order type at most $\omega^{n}$
@@
We remark that in our main results we are dealing with reflection iterated along arbitrary computable well-orders, while classically Feferman and Turing considered iterations along ordinal notations from Kleene’s $\mathcal{O}$. In fact the results are sensitive to this distinction and in Section 6 we establish $\omega^{n+1}+1$ as an upper bound on the order types of elements of $\mathcal{O}$ necessary to attain arbitrary true $\Pi_{2n+1}$-sentences as the consequences of iterated full uniform reflection.
@@ 1.2. Acknowledgements
The first author would like to thank Stella Moon whose questions lead him to start investigations on Feferman’s theorem. He would also like to thank Nikolay Bazhenov for pointing him to the work of Ash and Knight. Finally, we would like to thank Andrey Frolov for sharing with us his neat alternative proof of the theorem of Ash and Knight.
@@ 2.1. Computable orders
Usually recursive progressions progress proceed along some ordinal given by some ordinal notation system such as Kleene’s $\mathcal{O}$. We deviate from this since we will work directly with indices for computable orderings.
@@
Formally within first-order arithmetic we represent computable orders $\mathcal{L}$ as natural numbers coding pairs consisting of an index of a $\Pi_{1}$-set and an index of a $\Sigma_{1}$-set such that they represent the same set $\preceq^{\mathcal{L}}$ and this set is a binary relation satisfying the axioms of linear orders. We have first-order arithmetical formula $\mathsf{LO}(\mathcal{L})$ expressing that $\mathcal{L}$ is indeed a computable linear order in the sense above. We say that $\mathcal{L}$ like this is $\mathsf{PA}$-verifiable computable linear order if $\mathsf{PA}$ proves $\mathsf{LO}(\mathcal{L})$. Over $\mathsf{ACA}_{0}$ we formulate the second-order formula $\mathsf{WO}(\mathcal{L})$ expressing that $\mathcal{L}$ is a well-order:
@@
For the convenience of the reader we give a proof of Lemma 1 using standard techniques.
@@
where $\theta(y,x)$ is a quantifier-free formula in the language with primitive-recursive functions that expresses that
@@
The linearity of this $\mathcal{L}$ is of course trivial to verify in $\mathsf{ACA}_{0}$. And since $\mathsf{PA}$ is precisely the set of first-order consequences of $\mathsf{ACA}_{0}$, the produced order $\mathcal{L}$ is $\mathsf{PA}$-verifiable. ∎
@@ 2.2. Transfinite induction
Let $X$ be in the second-order part of $\mathfrak{A}$ such that $\varphi(x)$ defines $X$. Then, as $\mathfrak{A}$ satisfies induction for $\varphi(x)$, it also satisfies induction for $X$ and as $X$ was arbitrary, it satisfies set induction. If $\varphi$ is an arithmetical formula with parameters $Y_{1},\dots Y_{n}$ from $\mathfrak{A}$, then let $\theta_{i}$ be the arithmetical first-order formulas defining $Y_{i}$ and $\psi$ be the formula with every subformula of the form $t\in Y_{i}$ replaced by $\theta_{i}(t)$ for any term $t$. Then
@@
Given that we assumend $\mathcal{L}$ to be $\mathsf{PA}$-verifiable computable linear order, we conclude that we have $\mathfrak{A}\models\mathsf{ACA}_{0}+\mathsf{WO}(\mathcal{L})$. ∎
@@ 2.3. Uniform reflection and its iterations
Notice that $\mathsf{RFN}(\mathsf{RFN}^{(\mathcal{L},b)}(T))$ as a set of theorems might depend on the choice of $\Sigma_{1}$-formula recognizing axioms of $\mathsf{RFN}^{(\mathcal{L},b)}(T)$. Thus, in order to make the definition above formally correct we in fact need to produce a family of formulas defining axioms of theories $\mathsf{RFN}^{(\mathcal{L},a)}(T)$.
@@
Notice that the instance of $\mathsf{RFN}(\mathsf{RFN}^{(\mathcal{L},z)}(T))$ formulated as in Eq. 1 employs $\mathsf{Prv}_{\mathsf{RFN}^{(\mathcal{L},z)}(T)}(x)$ as a subformula, which in turn is fully determined by the formula $\mathsf{Ax}_{\mathsf{RFN}^{(\mathcal{L},z)}(T)}(x)$. For the purpose of the fixed point definition of $\mathsf{Ax}_{T}^{\mathcal{L}}(x,y)$, this instances of $\mathsf{Ax}_{\mathsf{RFN}^{(\mathcal{L},z)}(T)}(x)$ are as we intended simply $\mathsf{Ax}_{T}^{\mathcal{L}}(\underline{z},x)$. That is, above we define $\mathsf{Ax}_{T}^{\mathcal{L}}(x,y)$ in terms of its own Gödel number. Since the conjunction of the items (1) and (2) is clearly $\Sigma_{1}$, the standard proof of the Diagonal Lemma will produce in this case a $\Sigma_{1}$-formula $\mathsf{Ax}_{T}^{\mathcal{L}}(x,y)$.
@@
$\mathsf{RFN}^{\mathcal{L}}(T)=T+\bigcup\limits_{a\in\mathcal{L}}\mathsf{RFN}^{(\mathcal{L},a)}(T).$
@@
$\mathsf{RFN}^{(\mathcal{L},a)}(\mathsf{PA})\vdash\varphi$
@@ 3. Feferman’s Completeness Theorem via Deduction Chains
One might wonder whether Feferman could have used a different and technically more transparent approach other than his intricate applications of the recursion theore other than his intricate applications of the recursion theoremm. Indeed, the idea we are putting forward here is that two results available at the time could have led to such a proof. The first result, perhaps unsurprisingly, is Löb’s theorem from 1955 [lob1955solution] while the second is Schütte’s 1956 completeness theorem [schuette1956, Satz 6] for $\omega$-logic, which showed that with every true arithmetic statement a canonical primitive recursive proof tree (indeed a Kalmár-elementary tree) can be associated, entailing that $\omega$-logic with the primitive recursive $\omega$-rule (indeed Kalmár-elementary $\omega$-rule) is already complete. [^3] The completeness with the recursive $\omega$-rule is usually credited to Shoenfield whose article [shoenfield1959] appeared in 1959, while the canonical proof tree in many papers is associated with Mints’ work (see [mints1976], [sundholm1983], [franzen2004]). Löb’s theorem, of course, was well known at the time, but even today it remains largely unknown that Schütte was the first mathematician who showed the primitive recursive completeness of $\omega$-logic and introduced canonical search trees. [^4] The reasons for this are unclear. Schütte’s paper was in German, but that shouldn’t have been an obstacle back then. It appeared in the rather new journal Archiv für mathematische Logik und Grundlagenforschung. At the time, political reasons might have played a role in that there were two newly founded German journals of mathematical logic housed in different newly formed German states, perhaps appealing to different types of political allegiance.
@@
(The Tait calculus for $\omega$-logic) For classical logic, the so-called Tait calculus is very convenient (see [tait1968] and [schwichtenberg1977]). In it one deduces finite sequences of formulas, called sequents. Sequents are referred to by capital Greek letters $\Delta,\Gamma,\Lambda,\Xi,\ldots$. The result of appending a formula $\theta$ to a sequent $\Gamma$ is written $\Gamma,\theta$. Formulas are assumed to be in negation normal form, i.e., negations appear only in front of atomic formulas, whereas negation of more complex formulas is a defined operation, using De Morgan’s laws and dropping double negations. The other logical particles are ,,$\wedge$,$\vee$,$\exists$,$\forall$ , ∨ , ∃ , ∀. Literals are either atomic formulas or negated atomic formulas. Working in $\omega$-logic means that we can assume that formulas have no free variables. Below we are only concerned with $\omega$-logic based on the language of $\mathsf{PA}$. As a result, all literals are either true or false, and this can be checked by a primitive recursive algorithm. The axioms of the infinitary calculus are those sequents that contain a true literal. The rules are the usual ones for deducing a conjunction, disjunction, or existential formula, plus the $\omega$-rule (for details see [schwichtenberg1977]).
@@
Kleene’s normal form theorem for arithmetic formulas (Lemma 1) also assigns a linear computable ordering $\prec_{\varphi}$ to every sentence $\varphi$ such that the truth of $\varphi$ is equivalent to $\prec_{\varphi}$ being a wellordering. This result is obtained by a series of syntactic transformations performed on $\varphi$, and the feeling is that it doesn’t yield any new insights as to the truth of $\varphi$. Schütte’s completeness theorem, however, associates with $\varphi$ an ordering coming from a search tree for $\varphi$ in $\omega$-logic and the natural feeling is that this result is the more important one.

### v2

@@ 1. Introduction
By Gödel’s Second Incompleteness Theorem no consistent extension of $\mathsf{PA}$ with a computably enumerable axiomatization can prove its own consistency. Thus, whenever we extend a sound arithmetical theory $T$ to a sound theory $T^{\prime}$ that proves the consistency of $T$, it is guaranteed that $T^{\prime}$ is stronger than $T$.The idea of using this phenomenon to attain transfinite sequences of theories of ascending strength goes back to Alan Turing [turing1939]. In modern terms, Turing defined transfinite iterations of local reflection $\mathsf{Rfn}$ roughly as
@@
A subtle feature of this definition is that at each step of the inductive definition, one has to pick a c.e. axiomatization of $\mathsf{Rfn}^{\alpha}(T)$. Thus $\alpha$ has to be a computable ordinal with a fixed computable presentations, hence the theories $\mathsf{Rfn}^{\alpha}(T)$ depend not only on the order-type of $\alpha$ but also on the particular computable presentation. As was proven already by Turing, for any true $\Pi_{1}$-sentence $\varphi$ there is a computable presentation $\alpha$ of the ordinal $\omega+1$, in fact an ordinal notation in Kleene’s $\mathcal{O}$, such that $\mathsf{Rfn}^{\alpha}(\mathsf{PA})$ proves $\varphi$. Of course, there cannot be a single computable presentation $\alpha$ such that $\mathsf{Rfn}^{\alpha}(\mathsf{PA})$ proves every true $\Pi_{1}$ sentence, as this would mean that there is a consistent c.e. extension of $\mathsf{PA}$ proving all true $\Pi_{1}$ statements. Thus, the construction of theories $\mathsf{Rfn}^{\alpha}(\mathsf{PA})$ has an intensional character: Their consequences depend not only on the order type of $\alpha$ but also on the choice of a particular computable presentation.
Turing also conjectured that for every arithmetical theorem $\varphi$, there is a computable presentation of an ordinal $\alpha$ such that $\mathsf{Rfn}^{\alpha}(\mathsf{PA})$ proves $\varphi$. Later, Feferman [feferman1962] refuted this conjecture. However, he considered a variant of Turing’s progression, where local reflections $\mathsf{Rfn}(T)$ are replaced with uniform reflections $\mathsf{RFN}(T)$:
@@
For this variant, Feferman proved that Turing’s conjecture holds, i.e., any true arithmetical sentence is provable in $\mathsf{RFN}^{\alpha}(\mathsf{PA})$. This result is often referred to as Feferman’s completeness theorem. Feferman’s construction gives $\omega^{{\omega}^{\omega+1}}$ as an upper bound on the order type of the ordinal notations along which one needs to iterate reflection to obtain a proof of $\varphi$.
The presence of an upper bound in Feferman’s result indicates that it is a negative result: It essentially shows that proof-theoretic strength of an arithmetical sentence $\varphi$ couldn’t be appropriately measured by the least order-type of ordinal notation, iteration of reflection along which proves $\varphi$. However, when restricted to natural ordinal notation systems the length of iterations of reflection necessary to obtain a particular theory does reflect the proof-theoretic strength of the theory. This phenomenon was explored and used for the applications to ordinal analysis [schmerl1979fine, beklemishev2003proof]. We furthermore note that as was shown by the first author and James Walsh [pakhomov2021reflection], for systems of second-order arithmetic and their $\Pi^{1}_{1}$-consequences, the strength of the systems can be appropriately measured in terms of the length of iterations of $\Pi^{1}_{1}$-reflection, which is a measure closely connected with the $\Pi^{1}_{1}$-proof theoretic ordinals of theories.
Feferman’s original proof of his completeness theorem is not widely known, perhaps because of its considerable technical difficulty. One feature that Feferman’s proof shares with Turing’s (rather simple) proof of the $\Pi^{0}_{1}$-completeness of consistency progressions is the employment of non-standard definitions of the axioms of theories. These non-standard definitions are engineered so that we know them to be an axiomatization of a “natural” theory only if we know a certain sentence $\varphi$ to be true. In Feferman’s proof, they appear in ever more intertwined ways. Very artfully, he applies increasingly more entangled versions of the recursion theorem to generate non-standard definitions of theories (see [franzen2004] for further details). Franzén [franzen2004, p. 387] expressed this as follows:
The completeness theorem [i.e., Feferman’s] can be seen as a dramatic illustration of the role of intensionality in logic.
In this paper, we present three new proofs of Feferman’s completeness theorem. The first is a simple proof of this theorem using results not available to Feferman at the time such as the conservativity of $\mathsf{ACA_{0}}$ over $\mathsf{PA}$. The second proof is an alternative proof using techniques that Feferman could have had access to, such as Schütte’s completeness theorem for $\omega$-logic. At last, we combine our first proof with results of Ash and Knight [ash1990] in computable structure theory to obtain precise bounds on the order types appearing in Feferman’s theorem.
@@ 1.1. Structure of the paper
    For any computable linear order $\mathcal{L}$ such that $\mathsf{ACA}_{0}\vdash\mathsf{LO}(\mathcal{L})$ the theories $\mathsf{ACA}_{0}+\mathsf{WO}(\mathcal{L})$ and $\mathsf{PA}+\mathsf{TI}(\mathcal{L})$ prove exactly the same first-order sentences.
@@
The ideas of the proofs of these components are the following:
@@
    This is a particular case of $\mathsf{ACA}_{0}$-provable $\Pi^{1}_{1}$-completeness of well-orders. For the convenience of readers, we give a standard proof of the fact (Lemma 1).
@@
    We prove this in Lemma 17 by showing that the extension of any model of $\mathsf{PA}+\mathsf{LO}(\mathcal{L})+\mathsf{TI}(\mathcal{L})$ by the second-order universe consisting of first-order definable sets is a model of $\mathsf{ACA}_{0}+\mathsf{WO}(\mathcal{L})$.
@@
    Like most other facts about the iterations of reflection principles, this fact is straightforward to prove using Löb’s Theorem.
We note that neither of these facts is actually new. The first goes back to Kleene and the formalized version can be found in [simpson2009]. The second fact is a triviality, while the third one is folklore.
@@
To prove Theorem 1.1 we replace the use of $\Pi^{1}_{1}$-completeness of computable well-orders with a sharper theorem of Ash and Knight [ash1990], who proved that the set of indices of computable orderings of order-type $\omega^{n}$ is $\Pi^{0}_{2n+1}$-hard.
We complement Theorem 1.1 with the following result proving that the upper bound on order-types is in fact sharp:
@@
theoremlowerbound There exists a true $\Pi_{2n}$-sentence $\varphi$ such that for any computable $\mathsf{PA}$-verifiable linear order $\mathcal{L}$ of order-type at most $\omega^{n}$
@@
We remark that in our main results, we are dealing with reflection iterated along arbitrary computable well-orders, while classically Feferman and Turing considered iterations along ordinal notations from Kleene’s $\mathcal{O}$. In fact the results are sensitive to this distinction and in Section 6 we establish $\omega^{n+1}+1$ as an upper bound on the order types of elements of $\mathcal{O}$ necessary to attain arbitrary true $\Pi_{2n+1}$-sentences as the consequences of iterated full uniform reflection.
@@ 1.2. Acknowledgements
The first author would like to thank Stella Moon whose questions led him to start investigations on Feferman’s theorem. He would also like to thank Nikolay Bazhenov for pointing him to the work of Ash and Knight. We would like to thank Andrey Frolov for sharing with us his neat alternative proof of the theorem of Ash and Knight. Finally, we would like to thank the anonymous referee, as well as Mateusz Łełyk, Patryk Szlufik, and Patrick Lutz for their useful comments on an earlier version of the paper.
@@ 2.1. Computable orders
Usually, recursive progressions proceed along some ordinal given by some ordinal notation system such as Kleene’s $\mathcal{O}$. We deviate from this since we will work directly with indices for computable orderings.
@@
Formally within first-order arithmetic, we represent computable orders $\mathcal{L}$ as natural numbers coding pairs consisting of an index of a $\Pi_{1}$-set and an index of a $\Sigma_{1}$-set such that they represent the same set $\preceq^{\mathcal{L}}$ and this set is a binary relation satisfying the axioms of linear orders. There is a first-order arithmetical formula $\mathsf{LO}(\mathcal{L})$ expressing that $\mathcal{L}$ is indeed a computable linear order in the sense above. We say that an order $\mathcal{L}$ is a $\mathsf{PA}$-verifiable computable linear order if $\mathsf{PA}$ proves $\mathsf{LO}(\mathcal{L})$. Over $\mathsf{ACA}_{0}$ we formulate the second-order formula $\mathsf{WO}(\mathcal{L})$ expressing that $\mathcal{L}$ is a well-order:
@@
For the convenience of the reader, we give a proof of Lemma 1 using standard techniques.
@@
where $\theta(y)$ is a quantifier-free formula in the language with primitive-recursive functions that expresses that
@@
The linearity of this $\mathcal{L}$ is, of course, trivial to verify in $\mathsf{ACA}_{0}$. And since $\mathsf{PA}$ is precisely the set of first-order consequences of $\mathsf{ACA}_{0}$, the produced order $\mathcal{L}$ is $\mathsf{PA}$-verifiable. ∎
@@ 2.2. Transfinite induction
Let $X$ be in the second-order part of $\mathfrak{A}$ such that $\varphi(x)$ defines $X$. Then, since $\mathfrak{A}$ satisfies induction for $\varphi(x)$, it also satisfies induction for $X$ and since $X$ was arbitrary, it satisfies set induction. If $\varphi$ is an arithmetical formula with parameters $Y_{1},\dots Y_{n}$ from $\mathfrak{A}$, then let $\theta_{i}$ be the arithmetical first-order formulas defining $Y_{i}$ and $\psi$ be the formula with every subformula of the form $t\in Y_{i}$ replaced by $\theta_{i}(t)$ for any term $t$. Then
@@
Given that we assume $\mathcal{L}$ to be $\mathsf{PA}$-verifiable computable linear order, we conclude that we have $\mathfrak{A}\models\mathsf{ACA}_{0}+\mathsf{WO}(\mathcal{L})$. ∎
@@ 2.3. Uniform reflection and its iterations
Notice that $\mathsf{RFN}(\mathsf{RFN}^{(\mathcal{L},b)}(T))$ as a set of theorems might depend on the choice of $\Sigma_{1}$-formula recognizing axioms of $\mathsf{RFN}^{(\mathcal{L},b)}(T)$. Thus, in order to make the definition above formally correct, we, in fact, need to produce a family of formulas defining axioms of theories $\mathsf{RFN}^{(\mathcal{L},a)}(T)$.
@@
Notice that the instance of $\mathsf{RFN}(\mathsf{RFN}^{(\mathcal{L},z)}(T))$ formulated as in Eq. 1 employs $\mathsf{Prv}_{\mathsf{RFN}^{(\mathcal{L},z)}(T)}(x)$ as a subformula, which in turn is fully determined by the formula $\mathsf{Ax}_{\mathsf{RFN}^{(\mathcal{L},z)}(T)}(x)$. For the purpose of the fixed point definition of $\mathsf{Ax}_{T}^{\mathcal{L}}(x,y)$, this instance of $\mathsf{Ax}_{\mathsf{RFN}^{(\mathcal{L},z)}(T)}(x)$ is $\mathsf{Ax}_{T}^{\mathcal{L}}(\underline{z},x)$. That is, above we define $\mathsf{Ax}_{T}^{\mathcal{L}}(x,y)$ in terms of its own Gödel number. Since the conjunction of the items (1) and (2) is clearly $\Sigma_{1}$, the standard proof of the Diagonal Lemma will produce in this case a $\Sigma_{1}$-formula $\mathsf{Ax}_{T}^{\mathcal{L}}(x,y)$.
@@
$\mathsf{RFN}^{\mathcal{L}}(T)=T+\bigcup\limits_{a\in\mathcal{L}}\mathsf{RFN}(\mathsf{RFN}^{(\mathcal{L},a)}(T)).$
@@
$\mathsf{RFN}(\mathsf{RFN}^{(\mathcal{L},a)}(\mathsf{PA}))\vdash\varphi$
@@ 3. Feferman’s Completeness Theorem via Deduction Chains
One might wonder whether Feferman could have used a different and technically more transparent approach other than his intricate applications of the recursion theorem. Indeed, the idea we are putting forward here is that two results available at the time could have led to such a proof. The first result, perhaps unsurprisingly, is Löb’s theorem from 1955 [lob1955solution] while the second is Schütte’s 1956 completeness theorem [schuette1956, Satz 6] for $\omega$-logic, which showed that a canonical primitive recursive proof tree (indeed a Kalmár-elementary tree) can be associated with every true arithmetic statement, entailing that $\omega$-logic with the primitive recursive $\omega$-rule (indeed Kalmár-elementary $\omega$-rule) is already complete. [^2] The completeness with the recursive $\omega$-rule is usually credited to Shoenfield whose article [shoenfield1959] appeared in 1959, while the canonical proof tree in many papers is associated with Mints’ work (see [mints1976], [sundholm1983], [franzen2004]). Löb’s theorem, of course, was well known at the time, but even today it remains largely unknown that Schütte was the first mathematician who showed the primitive recursive completeness of $\omega$-logic [^3] This was confirmed by Göran Sundholm, an expert on the $\omega$-rule, in conversations with the second author. It is unclear why this is the case. Even though Schütte’s paper was in German that should not have been an obstacle back then. and introduced canonical search trees.
@@
(The Tait calculus for $\omega$-logic) For classical logic, the so-called Tait calculus is very convenient (see [tait1968] and [schwichtenberg1977]). In it, one deduces finite sequences of formulas, called sequents. Sequents are referred to by capital Greek letters $\Delta,\Gamma,\Lambda,\Xi,\ldots$. The result of appending a formula $\theta$ to a sequent $\Gamma$ is written $\Gamma,\theta$. Formulas are assumed to be in negation normal form, i.e., negations appear only in front of atomic formulas, whereas negation of more complex formulas is a defined operation, using De Morgan’s laws and dropping double negations. The other logical particles are ,,$\wedge$,$\vee$,$\exists$,$\forall$ , ∨ , ∃ , ∀. Literals are either atomic formulas or negated atomic formulas. Working in $\omega$-logic means that we can assume that formulas have no free variables. Below we are only concerned with $\omega$-logic based on the language of $\mathsf{PA}$. As a result, all literals are either true or false, and this can be checked by a primitive recursive algorithm. The axioms of the infinitary calculus are those sequents that contain a true literal. The rules are the usual ones for deducing a conjunction, disjunction, or existential formula, plus the $\omega$-rule (for details see [schwichtenberg1977]).
@@
Kleene’s normal form theorem for arithmetic formulas (Lemma 1) also assigns a linear computable ordering $\prec_{\varphi}$ to every sentence $\varphi$ such that the truth of $\varphi$ is equivalent to $\prec_{\varphi}$ being a wellordering. This result is obtained by a series of syntactic transformations performed on $\varphi$, and it seems that it doesn’t yield any new insights as to the truth of $\varphi$. Schütte’s completeness theorem, however, associates with $\varphi$ an ordering coming from a search tree for $\varphi$ in $\omega$-logic and it seems that this result is the more important one.